# Finslerian Wormholes in squared trace gravity

### v1

@@ Abstract
The idea of traversable wormholes, idealized channels between separate regions of space-time, has intrigued physicists since their theoretical construction in general relativity (GR) Morris ; Visser1995 . Within the standard GR context, these wormholes usually require the use of exotic matter that breaks the null energy condition (NEC) to stabilize geometry Hochberg1997 ; Visser1989 . Nevertheless, exotic matter is hypothetical, and its inconsistency with established physics has driven the study of alternative theories of gravity to create wormholes without contradicting the energy conditions Lobo2017 . Among all the modified theories padmanabhan2007 ; Durrer2008 ; Sotiriou:2008rp ; DeFelice2010 ; Nojiri:2010wj ; Capozziello2011 ; Clifton:2011jh ; Nojiri:2017ncd , $\mathfrak{f}(\mathds{R},\mathds{T})$ gravity has received significant interest, where $\mathds{R}$ is the Ricci scalar and $\mathds{T}$ represents the trace of the energy-momentum tensor.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\varpi_{2}=0.195$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=0.175$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++
$\varpi_{2}=0.155$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=0.12$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++
$\varpi_{2}=0.105$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=0.10$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++
$\varpi_{2}=-0.13$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=-0.15$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ ++
$\varpi_{2}=-0.165$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=-0.17$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ ++
$\varpi_{2}=-0.18$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=-0.2$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.215$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.22$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.24$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.25$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.255$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.26$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.265$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.28$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.295$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.305$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.315$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.335$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.34$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.355$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.365$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.37$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.38$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.39$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.43$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.435$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.44$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.445$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.45$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.455$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.465$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.58,~~~r>2.6$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.62,~~~r>2.2$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.655,~~r>1.9$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.685,~~r>1.8$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.73,~~~r>1.6$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.815,~~r>1.4$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.825,~~r>1.4$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.84,~~~r>1.3$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.845,~~r>1.3$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.855,~~r>1.3$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.86,~~~r>1.3$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.865,~~r>1.3$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.87,~~~r>1.3$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.905,~~r>1.2$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.915,~~r>1.2$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.925,~~r>1.2$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.94,~~~r>1.2$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.945,~~r>1.2$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.965,~~r>1.2$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.97,~~~r>1.2$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.975,~~r>1.2$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.985,~~r>1.1$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.995,~~r>1.1$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
@@
$\varpi_{2}=0.195$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=0.175$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++
$\varpi_{2}=0.155$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=0.12$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++
$\varpi_{2}=0.105$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=0.10$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++
$\varpi_{2}=-0.13$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=-0.15$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ ++
$\varpi_{2}=-0.165$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=-0.17$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ ++
$\varpi_{2}=-0.18$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ ++ $\varpi_{2}=-0.2$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.215$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.22$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.24$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.25$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.255$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.26$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.265$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.28$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.295$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.305$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.315$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.335$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.34$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.355$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.365$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.37$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.38$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.39$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.43$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.435$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.44$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.445$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.45$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.455$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.465$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.58$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.62$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.655$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.685$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.73$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.815$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.825$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.84$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.845$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.855$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.86$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.865$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.87$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.905$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.915$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.925$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.94$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.945$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.965$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.97$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.975$ $\xi>-1$ ++ $\checkmark$ $\checkmark$ -
$\varpi_{2}=-0.985$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ - $\varpi_{2}=-0.995$ $\xi<-1$ ++ $\checkmark$ $\checkmark$ -
@@
We demonstrated that both exponential and power-law shape functions produce geometrically regular wormholes, free of horizons and characterized by positive energy densities. Our analysis further showed that the exoticity integral $\oint(\rho+p_{r})dV$ remains positive in wide parameter ranges, confirming the physical viability of these solutions. Compared with purely Riemannian wormhole models Tripathy , we found that the inclusion of Finsler-Randers geometry broadens the admissible parameter space, especially for $\varpi_{2}$, and reduces the need for fine-tuned matter distributions.
@@
In future research, generalizing the framework to include other alternative theories of gravity, such as $f(Q,T)$ or teleparallel equivalents, could broaden the parameter space for physically realistic wormhole models. Finally, exploring the interplay between Finslerian anisotropies and dark energy components may offer new perspectives on unifying wormhole physics with large-scale cosmic acceleration.

### v2

@@ Abstract
The idea of traversable wormholes, idealized channels between separate regions of space-time, has intrigued physicists since their theoretical construction in general relativity (GR) Morris ; Visser1995 . Within the standard GR context, these wormholes usually require the use of exotic matter that breaks the null energy condition (NEC) to stabilize geometry Hochberg1997 ; Visser1989 . Nevertheless, exotic matter is hypothetical, and its inconsistency with established physics has driven the study of alternative theories of gravity to create wormholes without contradicting the energy conditions Lobo2017 . Among all the modified theories padmanabhan2007 ; Durrer2008 ; Sotiriou:2008rp ; DeFelice2010 ; Nojiri:2010wj ; Capozziello2011 ; Clifton:2011jh ; Nojiri:2017ncd ; De Falco , $\mathfrak{f}(\mathds{R},\mathds{T})$ gravity has received significant interest, where $\mathds{R}$ is the Ricci scalar and $\mathds{T}$ represents the trace of the energy-momentum tensor.
@@
Despite these advances, wormhole solutions in Riemannian $\mathfrak{f}(\mathds{R},\mathds{T})$ and related modified gravity theories often remain viable only within narrow regions of the equation-of-state parameter space, particularly when classical energy conditions are imposed. In many cases, NEC-satisfying configurations require fine-tuned matter distributions or are restricted to specific choices of shape functions. This limitation motivates the exploration of geometric generalizations beyond the Riemannian framework. In this context, Finsler–Randers geometry offers a natural extension in which space-time anisotropy arises intrinsically from the geometry itself, providing a promising avenue for enlarging the physically admissible parameter space of wormhole solutions.
@@
    $\begin{cases}\rho+p_{r}=1.33\rho\geq 0,\\ \rho+p_{t}=0.535\rho\geq 0,\end{cases}\Rightarrow~~\textrm{NEC is satisfied}$
@@
    $\begin{cases}\rho\geq 0,\\ \rho+p_{i}\geq 0~~\forall i=\{r,t\}.\end{cases}\Rightarrow~~\textrm{WEC is satisfied}$
@@
    $\begin{cases}\rho+p_{i}\geq 0~~\forall i=\{r,t\},\\ \rho+p_{r}+2p_{t}=(1+0.08-0.68)\rho=0.4\rho\geq 0.\end{cases}\Rightarrow~~\textrm{SEC is satisfied}$
@@
    $\begin{cases}\rho\geq|p_{r}|=0.33\rho,\\ \rho\geq|p_{t}|=0.465\rho.\end{cases}\Rightarrow~~\textrm{DEC is satisfied}$
@@
$\varpi_{2}=0.195$ $\xi<-1$ ++ Positive Positive ++ $\varpi_{2}=0.175$ $\xi<-1$ ++ Positive Positive ++
$\varpi_{2}=0.155$ $\xi<-1$ ++ Positive Positive ++ $\varpi_{2}=0.12$ $\xi<-1$ ++ Positive Positive ++
$\varpi_{2}=0.105$ $\xi>-1$ ++ Positive Positive ++ $\varpi_{2}=0.10$ $\xi<-1$ ++ Positive Positive ++
$\varpi_{2}=-0.13$ $\xi>-1$ ++ Positive Positive ++ $\varpi_{2}=-0.15$ $\xi>-1$ ++ Positive Positive ++
$\varpi_{2}=-0.165$ $\xi<-1$ ++ Positive Positive ++ $\varpi_{2}=-0.17$ $\xi>-1$ ++ Positive Positive ++
$\varpi_{2}=-0.18$ $\xi>-1$ ++ Positive Positive ++ $\varpi_{2}=-0.2$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.215$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.22$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.24$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.25$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.255$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.26$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.265$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.28$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.295$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.305$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.315$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.335$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.34$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.355$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.365$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.37$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.38$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.39$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.43$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.435$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.44$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.445$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.45$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.455$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.465$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.58,~~~r>2.6$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.62,~~~r>2.2$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.655,~~r>1.9$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.685,~~r>1.8$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.73,~~~r>1.6$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.815,~~r>1.4$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.825,~~r>1.4$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.84,~~~r>1.3$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.845,~~r>1.3$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.855,~~r>1.3$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.86,~~~r>1.3$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.865,~~r>1.3$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.87,~~~r>1.3$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.905,~~r>1.2$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.915,~~r>1.2$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.925,~~r>1.2$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.94,~~~r>1.2$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.945,~~r>1.2$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.965,~~r>1.2$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.97,~~~r>1.2$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.975,~~r>1.2$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.985,~~r>1.1$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.995,~~r>1.1$ $\xi>-1$ ++ Positive Positive -
@@
$\varpi_{2}=0.195$ $\xi<-1$ ++ Positive Positive ++ $\varpi_{2}=0.175$ $\xi<-1$ ++ Positive Positive ++
$\varpi_{2}=0.155$ $\xi<-1$ ++ Positive Positive ++ $\varpi_{2}=0.12$ $\xi<-1$ ++ Positive Positive ++
$\varpi_{2}=0.105$ $\xi>-1$ ++ Positive Positive ++ $\varpi_{2}=0.10$ $\xi<-1$ ++ Positive Positive ++
$\varpi_{2}=-0.13$ $\xi>-1$ ++ Positive Positive ++ $\varpi_{2}=-0.15$ $\xi>-1$ ++ Positive Positive ++
$\varpi_{2}=-0.165$ $\xi<-1$ ++ Positive Positive ++ $\varpi_{2}=-0.17$ $\xi>-1$ ++ Positive Positive ++
$\varpi_{2}=-0.18$ $\xi>-1$ ++ Positive Positive ++ $\varpi_{2}=-0.2$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.215$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.22$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.24$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.25$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.255$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.26$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.265$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.28$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.295$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.305$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.315$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.335$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.34$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.355$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.365$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.37$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.38$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.39$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.43$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.435$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.44$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.445$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.45$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.455$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.465$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.58$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.62$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.655$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.685$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.73$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.815$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.825$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.84$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.845$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.855$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.86$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.865$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.87$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.905$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.915$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.925$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.94$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.945$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.965$ $\xi<-1$ ++ Positive Positive -
$\varpi_{2}=-0.97$ $\xi>-1$ ++ Positive Positive - $\varpi_{2}=-0.975$ $\xi>-1$ ++ Positive Positive -
$\varpi_{2}=-0.985$ $\xi<-1$ ++ Positive Positive - $\varpi_{2}=-0.995$ $\xi<-1$ ++ Positive Positive -
@@
We demonstrated that both exponential and power-law shape functions produce geometrically regular wormholes, free of horizons and characterized by positive energy densities. Our analysis further showed that the exoticity integral $\oint(\rho+p_{r})dV$ remains positive over a wide range of parameters, confirming the physical viability of these solutions. Compared to earlier Riemannian $\mathfrak{f}(\mathds{R},\mathds{T})$ wormhole studies Tripathy , the present Finslerian construction admits NEC-satisfying solutions for significantly wider ranges of the tangential equation-of-state parameter $\varpi_{2}$, including domains that are inaccessible in the Riemannian limit.
@@
In future research, generalizing the framework to include other alternative theories of gravity, such as $\mathfrak{f}(\mathds{Q},\mathds{T})$ or teleparallel equivalents, could broaden the parameter space for physically realistic wormhole models. Finally, exploring the interplay between Finslerian anisotropies and dark energy components may offer new perspectives on unifying wormhole physics with large-scale cosmic acceleration.